OpenStudy (allyson):

Find an equation of a parraell line y=2x+5 passing through the points (5,5) y=?

OpenStudy (yrivers36):

y=2x-5 slope will be the same since its parallel plug in your points 5=2(5)+b y=mx+b 5=10+b subtract 10 from both side and your left with b=-5 your equation will be y=2x-5

OpenStudy (allyson):

thanks! (:

OpenStudy (yrivers36):

yw